7. Calculate the homicide rate per 100,000 people for each state and province in 1997 and 2008. Which state and which province had relatively higher homicide rates each year? Which country seems to have the higher homicide rate? Write a couple o sentences describing your results. Table 2. Homicides and Population Counts for Select States/Provinces, 1997-2008 1997 2008 State/province New Jersey Iowa Alabama Texas California Nova Scotia Quebec Ontario Manitoba British Columbia Homicides Population HomicidesPopulation 338 52 426 1,327 2,579 24 132 178 31 116 8,053,000 2,852,000 4,139,000 19,439,00 32,268,000 936,100 7,323,600 11,387,400 1,137,900 3,997,100 376 76 353 1,374 2,142 12 92 176 54 8,883,000 3,003,000 4,882,000 24,327,000 36,757,000 936,600 7,753,000 12,936,300 1,206,000 4,383,800 Source: Census Bureau &Statistics Canada, reported in Healey, Prus, & Lieflander (2016)
Answer #1
1997 2008
State/province Homicides (H) Population (P) Homicide rate per 1000 poplation=H*1000/P Homicides (H) Population (P) Homicide rate per 1000 poplation=H*1000/P Rate difference=1997-2008
New Jersey 338 8053000 0.0420 376 8883000 0.0423 -0.0004
Iowa 52 2852000 0.0182 76 3003000 0.0253 -0.0071
Alabama 426 4139000 0.1029 353 4882000 0.0723 0.0306
Texas 1327 19439000 0.0683 1374 24327000 0.0565 0.0118
California 2579 32268000 0.0799 2142 36757000 0.0583 0.0216
Nova Scotia 24 936100 0.0256 12 936600 0.0128 0.0128
Quebec 132 7323600 0.0180 92 7753000 0.0119 0.0062
Ontario 178 11387400 0.0156 176 12936300 0.0136 0.0020
Manitoba 31 1137900 0.0272 54 1206000 0.0448 -0.0175
British Columbia 116 3997100 0.0290 117 4383800 0.0267 0.0023

From the homicide rate per 1000 population between the year 1997 and 2008, we can conclude that the homicide rates are decreasing at most of the states/ provinces except New Jersey, Iowa, and Manitoba.